A study indicates that 80% of American teenagers have tattoos. You randomly sample 5 teenagers. Find the probability that at least 4 will have tattoos. Round your answer to 4 decimal places.
Answer #1

Solution

Given that ,

p = 0.80

q = 1 - p = 1 - 0.80 = 0.20

n = 5

Using binomial probability formula ,

P(X = x) = (n C x) * p x * (1 - p)n - x

P(X 4 ) = 1 - P( x <4)

= 1 - (5 C 0) * 0.80 0 * (0.20)5 - (5 C 1) * 0.80 1 * (0.20)4 - (5 C 2) * 0.80 2 * (0.20)3 - (5 C 3) * 0.80 3 * (0.20)2

=1-0.2627

probability=0.7373